### v1

@@ 4.2.3 CNOT Count
In addition to general gate optimization, targeting specific quantum gate types for reduction can significantly enhance circuit performance under certain conditions. In the NISQ era, the implementation of two-qubit entangling gates, such as CNOT gates, presents greater challenges compared to single-qubit gates. This is due to longer execution times (single-qubit gate: 60 ns, CNOT gate: 660 ns) and higher error rates (single-qubit gate $\approx 1\times 10^{-4}$, CNOT gate $\approx 1\times 10^{-2}$) [^2] Data source: IBM-Q superconducting quantum computer Kyoto. Consequently, minimizing the number of CNOT gates in circuit designs has emerged as a primary optimization objective, given that CNOT gates serve as representative two-qubit operations. This focused approach to gate reduction addresses the specific constraints and error profiles characteristic of current quantum hardware implementations.
@@
Numerous studies [16, 39, 100, 184, 132, 169, 133] have analyzed and optimized the theoretical number of CNOT gates in quantum circuits by refining decomposition methods. These efforts have progressively reduced the theoretical upper bounds from $\mathcal{O}(n^{3}4^{n})$ , as proposed in [16], to $\frac{23}{48}4^{n}-\frac{3}{2}2^{n}+\frac{4}{3}$ in [133]. Regarding CNOT gate depth, Jiang et al. [88] demonstrates that any $n$-qubit CNOT circuit can be optimized to $\mathcal{O}(\max\{\log n,\frac{n^{2}}{(n+m)\log(n+m)}\})$ depth using $m$ auxiliary qubits.A comprehensive discussion and summary of recent advancements related to the theoretical number of CNOT gates is presented in [81].
@@
Recent studies [176, 26, 166, 42, 196, 149, 112, 110, 161, 58, 139, 6] have focused on optimizing CNOT count in specific quantum circuits. Bravy et al. [26] identifies optimal Clifford circuits with minimal CNOT gates for up to 6 qubits, while Staudacher et al. [176] introduces methods to reduce CNOT count without compromising T gate optimization. Furthermore, optimizations tailored to variational quantum algorithms aim to enhance their performance in the NISQ era. Investigations such as [210, 201, 131, 181, 35, 68, 10, 177, 63, 117] focus on optimizing VQE circuits, whereas [215, 118, 119] concentrate on QAOA circuit optimization.
4.2.4 T Count and T Depth
@@
Unlike optimization goals in the NISQ era, fault-tolerant quantum computing (FTQC) implementation costs are not solely limited to hardware considerations, but also depend significantly on the temporal and spatial overheads of error-correcting codes. In FTQC models, Clifford group elements often have relatively simple transversal implementations across many error-correcting codes [214], whereas non-Clifford elements (such as T gates) are more complex to implement. For instance, in Steane codes [4] and surface codes [57], T gates require magic state distillation to implement [25]. The space-time overhead for surface codes is analyzed in [30], where Clifford gates incur a space-time cost of $\mathcal{O}(c_{T}d^{3})$ with a constant factor $c_{T}$ of unit order, while T gates require $\mathcal{O}(C_{T}d^{3})$ with $C_{T}\approx 160\sim 310$. Here, spatial overhead refers to the number of physical qubits needed to encode a logical qubit, and temporal overhead denotes the duration of the encoding protocol [30]. The implementation cost of T gates significantly exceeds that of Clifford gates by orders of magnitude [163, 30]. Consequently, optimizing the quantity and depth of non-Clifford gates (T gates) is of paramount importance in FTQC.
@@
Similar to reducing the overall gate count, directly minimizing the number of T gates is an intuitive approach. Several studies, including [207, 137, 95, 176, 75, 163, 9, 2], focus on optimizing T count. Abdessaied et al. [2] attempt to consolidate and reduce T count by reducing the number of H gates. Nam et al. [137] optimize T count for approximate Quantum Fourier Transform circuits. Kissinger et al. [95] and Heyfron et al. [75] employ the ZX diagram to reduce T count by extracting and consolidating non-Clifford phases. Amy et al. [9] transform the T count optimization problem into a minimum distance decoding problem for Reed-Muller codes, guaranteeing output circuits with optimal T count.
@@
Parallel to efforts in optimizing circuit depth, several studies [207, 8, 7] have focused on exploiting T gate parallelism to minimize T gate depth in quantum circuits. These investigations primarily examine the effect of parallelizable T gates on reducing overall circuit operation time. Notably, Amy et al. [7] proposes a polynomial-time optimization algorithm that, in most instances, guarantees the resultant circuit exhibits minimal T gate depth.
4.3 Quantum Logic Circuit Optimization Methods
@@
Based on the characteristics of different optimization methods, we summarize them in Table 3. The table lists the optimization targets and circuit representation methods for each optimization algorithm. It can be seen that gate circuits dominate, while phase polynomial representations are mainly used for T gate optimization. We also list some specific features, such as whether the algorithm requires ancillary qubits, whether it provides optimal solutions, and whether it optimizes parameterized circuits.
@@
Table 3: Review of quantum logical circuit optimization algorithms and characters.
@@
4.3.1 Brute Force Search
Brute force search algorithms find optimal circuits with specific depth and width through exhaustive enumeration. While this method guarantees optimal results, it incurs significant computational costs, limiting its application to small-scale logic circuits. Nevertheless, the optimal circuits generated by brute force search can be further utilized in other optimization algorithms. This section introduces the fundamental principles of brute force search algorithms.
@@ 4.3.1 Brute Force Search
The brute force search method typically comprises two core steps. First, a dataset containing all quantum logic circuits meeting specific criteria (e.g., certain depth and width) is constructed based on a set of candidate quantum gates. Subsequently, this dataset is traversed to identify the optimal circuit. As every subcircuit of an optimal quantum circuit is also optimal, the second step can employ bidirectional search or recursive backtracking to gradually identify the optimal quantum circuit composing the target unitary matrix from the dataset. However, these algorithms offer limited improvement in search efficiency. Therefore, investigating the optimization of dataset construction to reduce dataset volume and consequently decrease computational costs associated with searching is equally crucial.
@@
Amy et al. [8] utilize a breadth-first search algorithm to construct a gate circuit tree, where each edge corresponds to adding a gate from the candidate set. A child node’s circuit is derived by adding the gate on the edge to its parent node’s circuit. When searching for the optimal circuit of depth $k$, all nodes at depth $k$ in the tree are traversed. The study employs a bidirectional search method to reduce search depth, enabling the use of smaller datasets to search for larger circuits (using a dataset with maximum depth $k$ to construct all optimal quantum circuits with depth not exceeding $2k$). The correctness of bidirectional search stems from the fundamental fact that subcircuits of optimal circuits are also optimal. Another contemporary study, [99], proposes using unitary matrix equivalence classes to reduce dataset size. Unitary matrices that are equivalent under logical qubit reordering are grouped into the same class, significantly reducing dataset size (one equivalence class contains $n!$ unitary matrices). Bravyi et al.[26] investigate the construction of equivalence classes based on CNOT depth, considering nine different scenarios for adding CNOT gates (accounting for various placements of single-qubit H and P gates). Unitary matrices that are equivalent after removing single-qubit gates at both ends and reordering logical qubits are grouped into the same class, storing one representative element. For a target unitary matrix, reduction is first performed to find a representative element with the same CNOT depth, from which the optimal circuit is then derived. This methodology enables the search for all optimal circuits within six qubits.
The optimization methods employed in [199] and [153] are analogous to brute force search. They first construct an optimal circuit dataset of specified size, then reduce costs by replacing local segments of the target circuit with optimal ones. However, this approach only guarantees local optimality, not global optimality of the entire circuit. The algorithm proposed in [153] initially defines an $n$-qubit, $m$-layer grid, randomly populated with quantum gates from a set. Equivalence classes are then constructed based on unitary matrix equivalence within a certain precision. For the target circuit, all replaceable subcircuits are identified, their unitary matrices computed, and optimal replacements with minimal depth are sought from the dataset’s equivalence classes. Xu et al. [199] extends [153] from discrete to continuous quantum gate sets. The algorithm first constructs an optimal subcircuit dataset using breadth-first search and identifies equivalence classes for parameterized circuits. The research presents a verification method utilizing the function $|\bra{\psi_{0}}\mathbf{U}(\mathbf{p}_{0})\ket{\psi_{1}}|$ |, where $\mathbf{U}(\mathbf{p}_{0})$ represents a quantum circuit with parameters $\mathbf{p}_{0}$, |$\ket{\psi_{0}}$ and |$\ket{\psi_{1}}$ are random quantum states, and ||$\cdot$|| ⋅ | denotes complex modulus. By randomly sampling |$\ket{\psi_{0}}$, |$\ket{\psi_{1}}$, and $\mathbf{p_{0}}$, the function value of $\mathbf{U}$ can be calculated. If two unitary matrices yield identical results for one or more sets of variables, they are recognized as equivalent parameterized circuits. Utilizing this method, optimal replacement of certain local subcircuits within parameterized circuits can be implemented.
4.3.2 Pattern Matching
@@
Pattern matching is a widely employed technique in quantum circuit optimization, designed to identify and replace known patterns within a circuit to enhance its efficiency. Although this approach typically focuses on optimizing local sub-circuits and may not guarantee global circuit optimality, it offers advantages of low time complexity and favorable scalability. In quantum circuit optimization, pattern matching typically involves predefined equivalence substitution rules. The basic process of pattern matching algorithms is to set up a group of sub-circuit substitution rules in advance. These rules are then applied by traversing through sub-circuits using methods such as circuit partitioning or sliding windows. When applicable, these substitution rules are executed to reduce the cost of sub-circuits.
Peephole Optimization: Peephole optimization, a well-established technique in classical compilation optimization [127], focuses on small local segments such as a few instructions in classical computing or a limited number of quantum gates in quantum logic circuits. In the context of quantum circuit optimization, peephole optimization involves traversing the circuit to identify optimizable local patterns. We first introduce a special form of peephole optimization based on qubit states [112]. This method analyzes qubit states using a quantum state annotation (analogous to finite state machines in classical compilers). During circuit execution, optimization is applied when a qubit is found in a predefined pure or basis state and a local sub-circuit matches a predefined pattern. This algorithm is distinctive in its alteration of the optimized circuit’s unitary matrix while preserving functionality. Maslov et al. [121] and He et al. [69] optimize multi-control NOT gates by substituting Toffoli gates with phase Toffoli gates to reduce implementation costs. These two phase Toffoli gates differ from the standard Toffoli gate by a phase factor when acting on certain quantum states. However, due to their even times of applications, these phase factors cancel out in the final circuit. Consequently, the global effect is equivalent to that of the Toffoli gate, but with a reduced implementation cost. Sivarajah et al. [173] incorporates peephole optimization as a subroutine within the $\rm t\ket{ket}$ compiler, focusing on efficient optimization of specific quantum patterns. Bravyi et al. [27] introduces symbolic peephole optimization, addressing limitations in applying pattern matching optimizations where the replacement sub-circuit must be fully decoupled from other circuit components. This method decomposes CNOT gates into projection operators on control qubits and symbolic Pauli operators on target qubits. Symbolic Pauli operators include exponential term labels indicating control qubit indices, optimizing the decomposed circuits using dynamic programming, and predefined circuit equivalence rules before restoration.
@@
Recursive Decomposition: In addition to iterative pattern searching for optimization, recursive methods can also be employed to identify patterns within circuits. Recursive optimization is primarily applied to multi-controlled NOT gates. The core concept involves decomposing a controlled-NOT gate circuit with $c$ control qubits into circuits containing controlled-NOT gates with $c_{1}$ and $c_{2}$ control qubits, where $c_{1}$ and $c_{2}$ satisfy $c_{1}+c_{2}=c$. A common decomposition approach sets $c_{1}=c-1$ and $c_{2}=1$, effectively reducing one control qubit per iteration. In [14], a controlled-NOT gate with $c$ control qubits is recursively replaced by a controlled-NOT gate with $c-1$ control qubits and CNOT gates, utilizing quantum Karnaugh maps to verify circuit equivalence. Miller et al. [129] adopt a more generalized decomposition strategy, breaking down a controlled-NOT gate with $c$ control qubits into controlled-NOT gates with $c_{1}$ and $c_{2}$ control qubits. It then iterates through all possible $(c_{1},c_{2})$ value pairs, generating $c-2$ combinations and selecting the most cost-effective one. This recursive approach ultimately determines the optimal implementation.
@@
Divide-and-Conquer: Divide-and-conquer optimization first divides the complete circuits into sub-circuits of specified depth and width. These sub-circuit modules are then optimally reconstructed and subsequently merged. The QGo algorithm proposed in [196] focuses on optimizing CNOT gates (including SWAP gates) in quantum circuits with logical-to-physical qubit mapping. To partition an $n$-qubit circuit into multiple independent $k$-qubit circuit blocks (where gates in each block operate only on the corresponding $k$ qubits, with module depth not exceeding the complete circuit depth), QGo employs a greedy heuristic algorithm. It first constructs a DAG-based quantum gate dependency graph, where dependencies arise from the order of gate operations on qubits. Subsequently, it traverses all possible circuit partitions, scoring each using a CNOT count-based heuristic function. The highest-scoring partition is greedily selected, and its gates are removed from the current circuit. This process continues until the circuit is empty, signifying completed partitioning. QGo then employs a known optimal circuit synthesis algorithm [41] to reconstruct the decomposed circuit modules. Finally, all optimized circuits are merged, with potential gate eliminations at block interfaces. Prasad et al. [156] employs a similar divide-and-conquer approach to [196], but utilizes a linear, array-like data structure to describe quantum circuits, which can be viewed as the result of a depth-first search traversal of the DAG. Building upon [196],Patel et al. [149] proposes the QEst method, which replaces precise synthesis with approximate synthesis when composing optimal sub-circuits, achieving simpler circuit implementations within acceptable error margins.
Commutation: Commutation-based methods involve rearranging commutable gates to group specific gate types together, followed by other optimization techniques. Both [2] and [136] focus on optimizing the Clifford+T set. When optimizing T gates based on phase polynomials, H gates interfere with phase merging. Thus, these investigations aim to optimize T gates by reducing the number of H gates. Both approaches seek gates that commute with H for swapping, as consecutive H gates can cancel each other out. Clustering H gates together can reduce their overall count in the circuit, facilitating phase reduction. Due to the high computational complexity of determining commutativity between a quantum gate and a subcircuit, Nam et al. [136] propose a set of sufficient (but not necessary) rules to assess gate-subcircuit commutativity. The optimization method in [27] leverages gate commutation rules to relocate Pauli and SWAP gates, partitioning the circuit into three sections: Pauli gates, SWAP gates, and computational components (including CNOT, H, and S gates), allowing for tailored optimization strategies for each section. Diverging from direct quantum logic circuit optimization, Hietala et al. [76] introduce SQIR (Small Quantum Intermediate Representation), a characterization method between gate circuits and quantum assembly language. SQIR enables more efficient quantum program verification and equivalence checking. The VOQC optimization program proposed in this study utilizes a set of commutable gate exchange rules, combined with known single-qubit and two-qubit gate elimination rules, to optimize circuits represented in SQIR.
@@
Template Matching: Template matching is a specialized extension of pattern matching, wherein we consider a sequence of quantum gates that collectively result in the identity transformation (i.e., the unitary gate $\mathbf{I}$). In this context, the first half of this sequence can be represented by the inverse of the second half of the gates. The formal definition of template matching is as follows: suppose a template consists of $m$ quantum gates, which can be formulated as $\prod_{i=1}^{m}\mathbf{U}_{i}=\mathbf{I}$. Assume that in the quantum circuit yet to optimized, there exists a segment that matches the first $k\geq\left\lfloor\frac{m}{2}\right\rfloor+1$ gates of this template. Consequently, this segment can be replaced by:
$\mathbf{U}_{1}\mathbf{U}_{2}\cdots\mathbf{U}_{k}\rightarrow\mathbf{U}_{m}^{-1}\mathbf{U}_{m-1}^{-1}\cdots\mathbf{U}_{k+1}^{-1}.$ (13)
Given that $k\geq\left\lfloor\frac{m}{2}\right\rfloor+1$, the resulting circuit after substitution is guaranteed to be shorter than the original circuit.
Maslov et al. [122] employed template matching to optimize circuits containing NOT, CNOT, Toffoli, and MCT gates. They explored all templates of length less than or equal to $7$ and proposed strategies for optimization by traversing circuits bidirectionally to identify matching segments. Similar methodologies were applied in [124] with optimizations tailored to specific gate sets. They also categorized short templates satisfying additional conditions into specific circuit rules, such as gate commutativity. Additionally, Maslov et al. [123] introduced two template matching-based algorithms to optimize total gate count and circuit depth. The gate count optimization algorithm initiates with smaller templates, traverses the circuit from left to right, and identifies replaceable sub-circuits in either direction. The depth optimization leverages gate commutativity to enhance the parallelism of gates on different qubits, thereby reducing circuit depth. Bravyi et al. [27] also utilized gate commutativity and template matching for circuit optimization. They segregated CNOT, H, and S gates using commutativity before applying template matching to optimize this circuit segment. Iten et al. [85] employed canonical forms of circuits (representing them as DAGs) to identify small templates or maximal template matches in large circuits. They utilized bidirectional matching, considering whether nodes in the graph’s predecessor or successor DAG could match specific templates. While greedy matching suffices for optimal solutions in forward matching, introducing nodes in backward matching may induce errors in forward matching. Consequently, they initially constructed a gate selection tree for backward matching before searching for optimal solutions within this structure.
Graph Contraction: Graph contraction methods are primarily used in the optimization of circuits represented by ZX diagrams. These methods target specific patterns in a local region of the ZX diagram. Several representative patterns include: 1) Local Complementation: Merging elements with special angles (e.g., ±$\pm\frac{\pi}{2}$) into adjacent angles. 2) Pivoting: Combining multiples of 0 or $\pi$ through rotation into adjacent nodes. 3) Elimination and Fusion: Elimination of Empty Units and Fusion of Adjacent Units in Clifford Circuits. Kissinger et al.[95] focus on optimizing phases in ZX diagrams to reduce T count after circuit decomposition. For non-Clifford phases in the circuit, phase gadgets are introduced to represent phases as separate units apart from the non-phase circuit component. These gadgets are isolated from the non-phase part by H gates, and subsequently, the ZX diagram undergoes rule-based contraction optimization to derive an equivalent ZX diagram with optimized phases. Staudacher et al. [176] propose the first method to minimize CNOT count without compromising T count and depth optimization. The research introduces a heuristic scoring system for local complementations, rotation, and reduction rules, enabling random or greedy selection of optimization strategies based on post-decision scores until circuit complexity cannot be further reduced. To optimize the CNOT count, the scoring system directly correlates with the number of connecting lines in the diagram, allowing for adaption to different optimization objectives by modifying the scoring criteria. The study emphasizes that indiscriminate fusion of adjacent units may not always yield positive outcomes; conversely, dismantling units might facilitate the execution of new rules.
4.3.3 Reduction

### v2

@@ IV-B3 CNOT Count
In addition to general gate optimization, targeting specific quantum gate types for reduction can significantly enhance circuit performance under certain conditions. In the NISQ era, the implementation of two-qubit entangling gates, e.g. CNOT gates, presents greater challenges compared to single-qubit gates. This is due to longer execution times (single-qubit gate: 60 ns, CNOT gate: 660 ns) and higher error rates (single-qubit gate $\approx 1\times 10^{-4}$, CNOT gate $\approx 1\times 10^{-2}$) [^2] Data source: IBM-Q superconducting quantum computer Kyoto. Consequently, minimizing the number of CNOT gates in circuit designs has emerged as a primary optimization objective, given that CNOT gates serve as representative two-qubit operations. This focused approach to gate reduction addresses the specific constraints and error profiles characteristic of current quantum hardware implementations.
@@
Numerous studies [176, 177, 178, 91, 179, 180, 181, 182, 183] analyzed and optimized the theoretical number of CNOT gates in circuits by refining decomposition methods. These efforts have progressively reduced the theoretical upper bounds from $\mathcal{O}(n^{3}4^{n})$ in [176], to $\frac{23}{48}4^{n}-\frac{3}{2}2^{n}+\frac{4}{3}$ in [181]. Regarding CNOT gate depth, [184] demonstrated that any $n$-qubit CNOT circuit can be optimized to $\mathcal{O}(\max\{\log n,\frac{n^{2}}{(n+m)\log(n+m)}\})$ depth using $m$ auxiliary qubits. A comprehensive discussion of recent advancements related to the theoretical number of CNOT gates is presented in [9].
@@
Recent studies [163, 147, 28, 165, 157, 158, 151, 30, 31, 185, 186, 65] focused on optimizing CNOT count in specific quantum circuits. [147] identified optimal Clifford circuits with minimal CNOT gates for up to 6 qubits, while [163] introduced methods to reduce CNOT count without compromising T gate optimization. Furthermore, optimizations tailored to VQAs aiming to enhance the performance in the NISQ era with [187, 188, 189, 190, 191, 192, 193, 32, 18, 34] focused on optimizing VQE circuits, whereas [194, 112, 113] concentrated on QAOA circuit optimization.
IV-B4 T Count and T Depth
@@
Unlike optimization goals in the NISQ era, fault-tolerant quantum computing (FTQC) implementation costs are not solely limited to hardware considerations, but also depend significantly on the temporal and spatial overheads of error-correcting codes. In FTQC models, Clifford group elements often have relatively simple transversal implementations across many error-correcting codes [195], whereas non-Clifford elements (e.g. T gates) are more complex to implement. For instance, in Steane codes [196] and surface codes [197], T gates require magic state distillation to implement [198]. The space-time overhead for surface codes is analyzed in [199], where Clifford gates incur a space-time cost of $\mathcal{O}(c_{T}d^{3})$ with a constant factor $c_{T}$ of unit order, while T gates require $\mathcal{O}(C_{T}d^{3})$ with $C_{T}\approx 160\sim 310$. Here, spatial overhead refers to the number of physical qubits needed to encode a logical qubit, and temporal overhead denotes the duration of the encoding protocol [199]. The implementation cost of T gates significantly exceeds that of Clifford gates by orders of magnitude [67, 199]. Consequently, optimizing the quantity and depth of non-Clifford gates (T gates) is of paramount importance in FTQC.
@@
Similar to reducing the overall gate count, directly minimizing the number of T gates is an intuitive approach [161, 154, 164, 163, 26, 67, 27, 160][160] attempted to consolidate and reduce T count by reducing the number of H gates.  [154] optimized T count for approximate Quantum Fourier Transform circuits. [164, 26] employed the ZX diagram to reduce T count by extracting and consolidating non-Clifford phases. In [27], the T count optimization problem was transformed into a minimum distance decoding problem for Reed-Muller codes, guaranteeing optimality.
@@
However, as quantum hardware has evolved to support a greater degree of parallel execution of gates, the optimization objective has shifted from minimizing T count to minimizing the number of sequential T-gate layers, known as T depth. This metric is more crucial for reducing the overall circuit execution time. Accordingly, studies have focused on exploiting T-gate parallelism to minimize T depth in quantum circuits [161, 62, 13]. These investigations primarily examined the effect of parallelizable T gates on reducing overall circuit operation time. Notably, [13] proposed a polynomial-time optimization algorithm that, in most instances, guarantees the resultant circuit exhibits minimal T depth.
IV-C Quantum Logic Circuit Optimization Methods
@@
Based on the characteristics of different optimization methods, we summarize them in Table III. The table lists the optimization targets and circuit representation methods for each optimization algorithm. It can be seen that gate circuits dominate, while phase polynomial representations are mainly used for T gate optimization. We also list some specific features, e.g., whether the algorithm requires ancillary qubits, whether it provides optimal solutions, and whether it optimizes parameterized circuits.
@@
TABLE III: Review of quantum logical circuit optimization algorithms and characters.
@@
IV-C1 Brute Force Search (BFS)
@@
BFS finds optimal circuits with specific depth and width through exhaustive enumeration. While this method guarantees optimal results, it incurs significant computational costs, limiting its application to small-scale logic circuits. Nevertheless, the optimal circuits generated by BFS can be further utilized in other optimization algorithms. The BFS method typically comprises two core steps. First, a dataset containing all quantum logic circuits meeting specific criteria (e.g., certain depth and width) is constructed based on a set of candidate quantum gates. Subsequently, this dataset is traversed to identify the optimal circuit. As every subcircuit of an optimal quantum circuit is also optimal, the second step can employ bidirectional search or recursive backtracking to gradually identify the optimal quantum circuit composing the target unitary matrix from the dataset. However, these algorithms offer limited improvement in search efficiency. Therefore, investigating the optimization of dataset construction to reduce dataset volume and consequently decrease computational costs associated with searching is equally crucial.
@@
The work [62] adopted a breadth-first search scheme to construct a gate circuit tree, where each edge denotes the addition of a gate. The study used a bidirectional search scheme to reduce the search depth, enabling the discovery of larger circuits from smaller datasets (e.g., finding optimal circuits up to depth $2k$ using a dataset with depth $k$). The bidirectional approach relies on the principle that subcircuits of optimal circuits are also optimal. Another contemporary study [144] proposed using unitary matrix equivalence classes to reduce the dataset size. By grouping unitary matrices that are equivalent under logical qubit reordering, they significantly compressed the search space. Building on this idea, [147] investigated the construction of equivalence classes based on CNOT depth, considering nine different scenarios for adding CNOT gates (accounting for various placements of single-qubit H and P gates). It groups unitary matrices that are equivalent after removing single-qubit gates and reordering qubits, allowing the search for all optimal circuits within six qubits by first reducing a target unitary to its representative element.
[…]
IV-C2 Pattern Matching
@@
It is a popular technique in quantum circuit optimization that identifies and replaces known patterns within a circuit for efficiency. While it typically focuses on optimizing local sub-circuits and does not guarantee global circuit optimality, it is favored for its low time complexity and scalability. The process of pattern matching begins by establishing a set of sub-circuit substitution rules. These rules are applied by traversing sub-circuits, often using methods like circuit partitioning or sliding windows. When applicable, substitution rules are executed to minimize sub-circuit cost.
[…]
@@
Recursive Decomposition: Recursive methods were often used to identify patterns in quantum circuits, particularly for multi-controlled NOT gates. These methods broke down a controlled-NOT gate circuit with $c$ control qubits into circuits with $c_{1}$ and $c_{2}$ control qubits, where $c=c_{1}+c_{2}$. A common decomposition approach sets $c_{1}=c-1$ and $c_{2}=1$, reducing one control qubit per iteration. In [155], this method replaced a controlled-NOT gate with $c$ control qubits by one with c1c-1 qubits and additional CNOT gates, verified using quantum Karnaugh maps. The approach in [53] generalized this strategy, considering all possible pairings of $(c_{1},c_{2})$, generating multiple combinations and selecting the most efficient one.
@@
Divide-and-Conquer: it splits circuits into sub-circuits of specified depth and width, which are optimized independently before being recombined. The QGo algorithm [157] optimized CNOT and SWAP gates by partitioning an $n$-qubit circuit into multiple $k$-qubit blocks. It employed a greedy heuristic based on a quantum gate dependency graph, selecting the best partitions and iterating until the entire circuit was decomposed. QGo then used optimal circuit synthesis [200] to reconstruct the sub-circuits. A similar approach in [156] employed a linear, array-like data structure to describe quantum circuits, which can be viewed as the result of a depth-first search traversal of the DAG. Building upon [157], QEst [158] replaced exact synthesis with approximate synthesis, achieving simpler circuit implementations with acceptable error margins.
Commutation: Commutation-based methods optimized circuits by rearranging commutable gates. The works in [160] and [63] focused on optimizing the Clifford+T set by reducing the number of T gates through phase polynomials. H gates, which interfered with phase merging, were rearranged to cancel out and minimize their count. [150] utilized commutation rules to optimize Pauli and SWAP gates by partitioning the circuit into three sections: Pauli gates, SWAP gates, and computational components (including CNOT, H, and S gates), optimizing each with tailored strategies. Additionally, [159] introduced SQIR (Small Quantum Intermediate Representation) for efficient quantum program verification and equivalence checking, combining commutation with gate elimination.
@@
Template Matching: it extends pattern matching by optimizing sequences of gates that result in the identity transformation, where the first half of the sequence is the inverse of the second. For a template consists of $m$ gates that satisfy $\prod_{i=1}^{m}\mathbf{U}_{i}=\mathbf{I}$, if a segment in the circuit matches the first $k$ gates of the template, it can be replaced by the inverse of the remaining gates $\mathbf{U}_{1}\mathbf{U}_{2}\cdots\mathbf{U}_{k}\rightarrow\mathbf{U}_{m}^{-1}\mathbf{U}_{m-1}^{-1}\cdots\mathbf{U}_{k+1}^{-1}$. The optimization reduces circuit length as long as $k\geq\left\lfloor\frac{m}{2}\right\rfloor+1$.
[54] employed template matching to optimize circuits containing NOT, CNOT, Toffoli, and MCT gates. They explored all templates of length $\leq 7$ and proposed strategies for optimization by traversing circuits bidirectionally to identify matching segments. Similar methodologies were applied in [55] with optimizations tailored to specific gate sets, and explored additional circuit rules, e.g. gate commutativity. Template matching algorithms can also optimize gate count and circuit depth [12], leveraging gate commutativity to enhance parallelism of gates. [150], meanwhile, employed commutation and template matching to optimize CNOT, H, and S gates. [150] introduced DAG-based template matching with templates as subgraphs. A bidirectional search finds the match for both predecessor and successor nodes in the DAG.
Graph Contraction: Primarily optimized circuits represented by ZX diagrams, targeting specific local patterns. Several representative patterns include: 1) Local Complementation: Merging elements with special angles (e.g., $\pm\frac{\pi}{2}$) into adjacent angles. 2) Pivoting: Combining multiples of 0 or $\pi$ through rotation into adjacent nodes. 3) Elimination and Fusion: Elimination of Empty Units and Fusion of Adjacent Units in Clifford Circuits.
In [164], phase optimization in ZX diagrams reduced T count after circuit decomposition. For non-Clifford phases, phase gadgets were introduced and isolated using H gates. Rule-based contraction optimization are then conducted to reduce the T count. Similarly, [163] minimized CNOT count without compromising T count by using a heuristic scoring system, allowing random or greedy selection of optimization strategies. The scoring system can be adapted to different objectives by modifying the criteria. The study emphasizes that indiscriminate fusion of adjacent units may not always yield positive outcomes; conversely, dismantling units might facilitate the execution of new rules.
IV-C3 Reduction
An alternative approach to addressing optimization problems is to reduce them to other problems and apply specialized solvers. A common method utilizes phase polynomials to represent quantum circuits, which are then converted into Reed-Muller codes. This transformation reduces the circuit optimization problem to a Reed-Muller decoding problem, enabling the use of dedicated solving techniques.